### v5

@@ 9 Conclusions: Why Nature Does Not Like Zombies
We have described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [62].
As previous research pointed out [36], the computer metaphor with hardware and software is a simplification. Rather, software is a state of hardware. Nature privileges efficiency over abstract simplification. Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they are adaptive at a every level [28, 43]. Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
One consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness. We have shown a consistent definition of access consciousness requires 1ST and 2ND order selves. Phenomenal consciousness arises first, access comes later. Unlike panpsychism, we don’t believe rocks are conscious. Only self-organising systems that need to adapt, motivated by valence, while keeping track of the self are conscious. Consciousness is an adaptation, and a philosophical zombie is impossible.
Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience. Carruth argued that if qualia are dispositional, then that is impossible [96]. Our arguments align with Carruth’s. A causal identity is inherently dispositional, as is embodiment.
In summary, we have argued that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world and others and iii) the self as modelled by others. These follow from learning and enacting a lattice of causal identities, if the scale and incentive preconditions are met. There is something it is like to be an organism with a self, and that is where phenomenal consciousness begins. Our proposal lays the foundation of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to the organic facts of humanity than the conceptual fiction of zombies.
Funding

### v6

@@ 9 Conclusions: Why Nature Does Not Like Zombies
In this paper we have proposed a novel, causal identity theory of consciousness, taking as a starting point not the conscious mind but the embodied experiencing organism. Building on Stack Theory Bennett (2025a), we have illustrated in formal terms how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as healthy humans can Pearl and Mackenzie (2018).
Box 4: Experimental results at a glance We report three simple Monte Carlo benchmarks from the associated workshop paper Bennett (2026b). They are minimal tests of the formal claims in noisy randomised settings. Experiment 1. Weakness beats simplicity. Across randomised abstraction layers, w-maxing has zero regret under a natural ignorance prior. A shortest description length selector incurs large regret. In one benchmark, that regret is about $13.4$ bits. That is an odds penalty of about $2^{13.4}\approx 10{,}800$. In other words description length depends on the encoding, while weakness is an invariant of the embodied language. Experiment 2. Probing defeats decoder mismatch. When a listener’s decoding rule is hidden, passive signalling stays near chance. With two short probes, a probing policy reaches about $0.93$ success across both stationary and nonstationary decoder settings. An oracle decoder reaches about $0.96$ in the stationary setting. In other words, causal intervention helps ascertain the listener’s decoder for communication. Experiment 3. Binding creates trust. In randomised trust games, nonbinding promises do not generate trust in equilibrium. Binding moves do. In one benchmark sample, binding yields trust about $0.95$ by horizon $H=5$. Third order selves thus help navigate complex social environments. Protocols and plots. See ESM Experiments S1 to S3, with Tables S1 to S4 and Figures S1 to S4.
One consequence of our approach is that it places the qualitative aspect of conscious embodied experience [^27] Inherently valenced information processing through the lens of self-preservation. before access consciousness. Ontologically, this means that before one can access information, one needs to experience information through bodily and causal interaction with the world. We have shown a consistent definition of access consciousness requires 1ST and 2ND order selves. Phenomenal consciousness arises first, access comes later. Unlike panpsychism, we don’t believe rocks are conscious. Only self-organising systems that need to adapt, motivated by valence, while keeping track of self-relevant information are conscious. Consciousness is a key adaptation of a vulnerable organism in a physical world.
In summary, we have argued that access consciousness at the human level is impossible without the ability to hierarchically model the following.
  1. 1.
    The self.
  2. 2.
    The world and others.
  3. 3.
    The self as modelled by others.
These follow from learning and enacting a lattice of causal identities, if the scale and incentive preconditions are met. There is something it is like to be an organism with a self. Paradoxically, phenomenal consciousness begins in organisms that have an end and a primary goal to avoid that end. Our proposal lays the foundation of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to the facts of embodiment than the fiction of zombies.
Acknowledgements